# abEunwoo Heo    abByeongchan Choi    abJae-Hun Jung Corresponding author

### v1

@@ 3.2 Toy example of path-representable distance
Consider a connected weighted graph $G=(V,E,W_{E})$ with $V=\{v_{1},v_{2},v_{3},v_{4}\}$ and $E=$$\{(v_{1},v_{2}),$ $(v_{2},v_{3})$, $(v_{3},v_{4})$,$(v_{1},v_{4})\}$ as shown in the left figure of Figure 3. How many path-representable distances can we make for any path choice function $g$ on $G$? A maximal path can represent any path choice function $g$ on $G$, where the maximal path refers to a path that is maximal with respect to the inclusion. The right figure of Figure 3 represents the maximal paths indicated by the given path choice function on graph $G$. The maximal paths are well-defined due to the consistency property of the path choice function. In this case, a total of eight possible path choice functions can be defined on $G$, as shown in Figure 3.
@@
For the connected weighted graph $G=(V,E,W_{E})$ shown in the left figure of Figure 4, all path-representable distances derived from all the path choice functions are illustrated in the right figure of Figure 4, with a total of eight possibilities. In this case, all the distances are weight-dominated, while only the distances $d_{5},d_{6},d_{7},d_{8}$ are cost-dominated.
@@ 3.3 Main theorem
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider two cost-dominated path-representable distances $d_{i}$ and $d_{j}$ by path choice functions $g_{i}$ and $g_{j}$ on $G$, respectively. If $d_{i}(v,w)\leq d_{j}(v,w)$ for any vertices $v$ and $w$, then there exists an injective function $\varphi_{i,j}:\textbf{bcd}_{1}(d_{i})\rightarrow\textbf{bcd}_{1}(d_{j})$ defined as $\varphi_{i,j}([a,b])=[a,c]$ such that $b\leq c$ for any $[a,b]\in\textbf{bcd}_{1}(d_{i})$.
Figure 5: The injective functions $\varphi_{i,j}:\textbf{bcd}_{1}(d_{i})\rightarrow\textbf{bcd}_{1}(d_{j})$, as described in Theorem 3.4, for the four cost-dominated path-representable distances $d_{5},d_{6},d_{7},d_{8}$ in Figure 4.
Figure 5 illustrates the application of Theorem 3.4 to the cost-dominated path-representable distances $d_{5},d_{6},d_{7},d_{8}$ described in Figure 4. The left figure of Figure 5 shows the partial order relationships among distances $d_{5},d_{6},d_{7},d_{8}$, while the right figure displays the injective functions $\varphi_{5,6},\varphi_{6,7},\varphi_{5,8},\varphi_{8,7}$ between their 1-dimensional persistence barcodes.
@@ 3.4 Remarks
The main theorem can also be formulated using the concept of a poset. Let $G$ be a connected weighted graph. To analyze the graph by defining the distance between the nodes, it is natural to define the set whose elements are the distances defined on $G$. Especially in our theory, we focus on the poset $(\mathcal{L},\leq)$ whose elements are the cost-dominated path-representable distances defined on $G$, where for two elements $d_{1},d_{2}\in\mathcal{L}$, $d_{1}\leq d_{2}$ if $d_{1}(v,w)\leq d_{2}(v,w)$ for any two points $v,w\in V$. Furthermore, we calculate persistent homology using the standard algorithm [4], which calculates persistence barcode from the given distance matrix. Note that $(G,d)$ gives a unique distance matrix. We define such a function as $\textbf{bcd}_{k}$, that is, $\textbf{bcd}_{k}(d)$ is a $k$-dimensional persistence barcode obtained by $(G,d)$. Then the codomain of the function will be the poset of persistence barcodes on $G$, as a subset of the poset of interval sets paired with $k$-simplices. Here, the partial order holds if we can find the proper inclusion between the elements with the same associated simplex in the sense of the main theorem. To be precise, for two $k$-dimensional persistence barcode $S_{1}$ and $S_{2}$, we define $S_{1}\leq S_{2}$ if there is an injection $\varphi:S_{1}\hookrightarrow S_{2}$ such that $\varphi([a,b]_{\sigma})=[a,b^{\prime}]_{\sigma}\text{ for any }[a,b]_{\sigma}\in S_{1}\text{ with }b\leq b^{\prime}.$ The main theorem claims that $\textbf{bcd}_{1}$ becomes the order-preserving map.
@@
First, if we choose $d_{weight}$ as an element, then it becomes the least element according to the Definition 2.1. Note that $d_{weight}$ is indeed the element in the domain by Proposition 3.3.
Now we take a counterexample for the greatest element. Let $G$ be a cycle graph with at least 5 vertices and suppose that $d_{\max}$ is the greatest element in $P$. As illustrated in Figure 6, fix a control point $v_{0}$. Since $d_{\max}$ is path-representable, for every other vertex $v$ of $G$, we must choose between two paths from $v_{0}$ to $v$ which $d_{\max}$ calculates. Let $v_{i}$ and $v_{i+1}$ be two consecutive nodes. Let $d_{1}$ be the distance that chooses the path from $v_{0}$ to $v_{i}$ that does not contain $v_{i+1}$ and the path from $v_{0}$ to $v_{i+1}$ that contains $v_{i}$. Let $d_{2}$ be the distance that chooses the path from $v_{0}$ to $v_{i+1}$ that does not contain $v_{i}$ and the path from $v_{0}$ to $v_{i}$ that contains $v_{i+1}$. Note that we can make these two distances cost-dominated since $v_{i}$ and $v_{i+1}$ are not adjacent to $v_{0}$, and $G$ is a cycle graph. Then $d_{1}(v_{0},v_{i})>d_{2}(v_{0},v_{i})$ and $d_{1}(v_{0},v_{i+1})<d_{2}(v_{0},v_{i+1})$. Thus, due to the assumption for $d_{\max}$, we have $d_{\max}(v_{0},v_{i})=d_{1}(v_{0},v_{i})$ and $d_{\max}(v_{0},v_{i+1})=d_{2}(v_{0},v_{i+1})$ since there are only two candidates for each selection of path. This leads to a contradiction to the consistency condition of path-representable distance.
Figure 7: $\textbf{bcd}_{1}$ being an order preserving map between the poset of cost-dominated path-representable distance on $G$ (Left) and the poset of $1$-dimensional persistence barcode on $G$ from Figure 4 (Right). Here the arrow denotes the partial order for each poset.
@@ 4 Proofs of main theory
For a weighted graph $G=(V,E,W_{E})$, let $d:V\times V\rightarrow\mathbb{R}$ be any distance defined on $G$. If $\sigma=(v,w)$ be a birth simplex of dimension $1$ with the distance $d(v,w)$ of a $1$-dimensional cycle, there is no vertex $v_{0}\in V(G)$ such that $d(v,v_{0})<d(v,w)$ and $d(w,v_{0})<d(v,w)$.
@@
Let $e=(v,w)$ be a birth simplex of dimension $1$ with a birth time $d(v,w)$. Suppose $e\notin E$. There is a path $g(v,w)$ in $G$ such that $d(v,w)=\sum_{e\in g(v,w)}W_{E}(e)$. Denote the path $g(v,w)$ as the sequence of vertices $v=v_{0},v_{1},v_{2},\ldots,v_{m-1},v_{m}=w$ and $a_{k}=W_{E}((v_{k-1},v_{k}))$. Then, by the consistency property of the path choice function $g$, the path $g(v,w_{m-1})$ is a sequence of $v_{0}$, $v_{1}$, $v_{2}$, $\ldots$, $v_{m-1}$. Therefore, we know that $d(v,v_{m-1})=a_{1}+a_{2}+\cdots+a_{m-1}<d(v,w)$. Since $d(v_{m-1},w)<d(v,w)$ is obvious, this contradicts Lemma 4.1. ∎
Let $\mathcal{B}$ be the set of all birth edges in $\textbf{bcd}_{1}(d)$. By Proposition 4.2, we know that $\mathcal{B}\subseteq E$. Generally, for any edge $e=(v,w)\in E$, it is not necessary that $W_{E}(e)=d(v,w)$. For example, cases where $W(e)>d_{\text{weight}}(v,w)$ are common. However, for the birth edges, the following interesting proposition holds.
@@
Let $G=(V,E,W_{E})$ be any connected weighted graph and $\mathcal{T}\in\text{MST}(G)$. Suppose that we have a cycle $\tau$ in $G$ and that there is an edge $e\in\tau$ such that $W_{E}(e)>W_{E}(e^{\prime})$ for any other edge $e^{\prime}\in E$. Then $e\notin\mathcal{T}$.
@@
Consider a connected weighted graph $G=(V,E,W_{E})$. Define $\bar{E}=\{(v,w)\mid v,w\in V\}$. For any path choice function $g$, we have its path-representable distance $d_{g}$ and a weighted complete graph $K^{g}=(V,\bar{E},W_{\bar{E}})$, where $W_{\bar{E}}(e)=d_{g}(v,w)$ for all $e=(v,w)\in\bar{E}$. We name this graph $K^{g}$ the the graph completion of $G$ by $g$.
@@
(iii) Consider a minimum spanning tree $\mathcal{T}^{\prime}\in\text{MST}(G)$. If $\mathcal{T}=\mathcal{T}^{\prime}$, then the proposition holds. Suppose that $\mathcal{T}\neq\mathcal{T}^{\prime}$, which means that there is an edge $e\in\mathcal{T}\setminus\mathcal{T}^{\prime}$ and a unique path $p^{\prime}$ in $\mathcal{T}^{\prime}$ connecting the endpoints of $e$. In the path $p^{\prime}$, there is an edge $e^{\prime}\in\mathcal{T}^{\prime}\setminus\mathcal{T}$ because $e+p^{\prime}$ is a cycle in $G$ and $\mathcal{T}$ is a tree. By $(ii)$, we have $W_{E}(e)=W_{\bar{E}}(e)$. Suppose that $W_{E}(e^{\prime})\neq W_{\bar{E}}(e^{\prime})$ and denote $e^{\prime}=(v,w)$. Then there is a path $g(v,w)$ such that $W_{E}(e^{\prime})>W_{E}(e_{i})$ for any $e_{i}\in g(v,w)$. This contradicts Lemma 4.4 because $e^{\prime}\in\mathcal{T}^{\prime}\in\text{MST}(G)$ so that $W_{E}(e^{\prime})=W_{\bar{E}}(e^{\prime})$. Since $\mathcal{T}^{\prime}\in\text{MST}(G)$, we have $W_{E}(e^{\prime})\leq W_{E}(e)$. Otherwise, by replacing $e^{\prime}$ with $e$ in the edge swap process, we have a contradiction as $\mathcal{T}^{\prime}\in\text{MST}(G)$. Similarly, since $\mathcal{T}\in\text{MST}(K^{g})$, we have $W_{\bar{E}}(e^{\prime})\geq W_{\bar{E}}(e)$. To summarize, the followings hold:
@@
Proposition 4.6 implies that any weight-dominated path choice function $g$ always keeps $\text{MST}(G)$ invariant, for which the weight-dominated condition is essential. Otherwise, neither $\text{MST}(G)$ nor $\text{MST}(K^{g})$ can be guaranteed to be larger. A counterexample is shown in Figure 8 below.
@@
For the fixed order $\pi$ on the edge set $\bar{E}$ in $K^{g}$, consider the set of all $1$-simplices $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$ for some 00-simplex $\sigma_{i}$, denoted by $\mathcal{T}^{\pi}$. Then
$\textnormal{MST}(K^{g})=\{\mathcal{T}^{\pi}\mid\pi\text{ is an order on edge set $\bar{E}$ in $K^{g}$}\}$
@@
()superset-of-or-equals($\supseteq$)( ⊇ ) The process of sequentially constructing a minimum spanning tree matches precisely the process of incrementally calculating $\text{low}_{\mathbf{R}}$. To explain in more detail, for every vertex in $K^{g}$, a 00-simplices (vertices) is connected to the $1$-simplices (edges) causing its death. According to the definition of $\text{low}_{\mathbf{R}}$, no cycles can form hence it is a spanning tree. Since $\text{low}_{\mathbf{R}}$ selects edges in order from the lowest value, it becomes the minimum spanning tree. Therefore, for any order $\pi$ on the edge set $\bar{E}$ in $K^{g}$, $\mathcal{T}^{\pi}$ is an element in $\text{MST}(K^{g})$.
()($\subseteq$)( ⊆ ) Consider an arbitrary element $\mathcal{T}\in\text{MST}(K^{g})$. Let $\pi$ be an order on $\bar{E}$ such that, for any given weights on the edges, if the weights are the same, the edges in $\mathcal{T}$ are preferred over those in $K^{g}\setminus\mathcal{T}$. When computing the persistent homology of $K^{g}$, even if the order of the $1$-simplices (edges) corresponding to the indices of the matrix in the $\mathbf{R}=\mathbf{D}\mathbf{V}$ algorithm is rearranged according to $\pi$, the result of the persistent homology computation remains unchanged (permutation invariant). In this case, $\mathcal{T}=\mathcal{T}^{\pi}$. ∎
@@
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider any weight-dominated choice functions $g_{1}$ and $g_{2}$ and theirs completions $K^{g_{1}}$ and $K^{g_{2}}$. Then there exists an order $\pi$ in $\bar{E}$ such that their minimum spanning trees on $K^{g_{1}}$ and $K^{g_{2}}$ induced by the standard algorithm are exactly the same.
@@
(i) First, we show that the $e$-column of the reduced $R_{1}$ is zero. From $e\in\mathcal{B}_{2}$, we know that the $e$-column of $\mathbf{R}_{2}$ is zero. The fact that the $e$-column of $\mathbf{R}_{2}$ is a zero column means that there are column indices $j_{1}$, $\ldots$, $j_{m}$ corresponding to edges $e_{j_{1}},\ldots,e_{j_{m}}$ such that $e+\sum_{l=1}^{m}e_{j_{l}}$ is a cycle and there are 00-simplices $\sigma_{i_{j_{k}}}$ such that $i_{j_{k}}=\text{low}_{\mathbf{R}_{2}}(j_{k})$. By Proposition 4.8, we can assume that the minimum spanning trees $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ for $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$ in the standard algorithm are exactly the same for some fixed order $\pi$ on edge set $\bar{E}$. Therefore, we know that $i_{j_{k}}=\text{low}_{\mathbf{R}_{1}}(j_{k})$ for all $k$. Thus the $e$-column of $\mathbf{R}_{1}$ is zero due to the column indices $j_{1},\ldots,j_{m}$ during the column operations in the standard algorithm.
(ii) Next, we show that if the $1$-cycle $\tau$ corresponding to the birth edge $e=\sigma_{i}$ is terminated by the $2$-simplex $\sigma_{j}$ in $\mathbf{R}_{1}$, then $f_{1}(\sigma_{i})<f_{1}(\sigma_{j})$. Consider a pair of simplices $(\sigma_{i},\sigma_{j})$ which is $\text{low}_{\mathbf{R}_{1}}(j)=i$. For any $t\in\mathbb{R}$, define $\mathbb{X}^{2}_{t}={f_{2}}^{-1}((-\infty,t])$ and $\mathbb{X}^{1}_{t}={f_{1}}^{-1}((-\infty,t])$. As $f_{2}\leq f_{1}$ on $\mathbb{X}$, we know that $\mathbb{X}^{1}_{t}\subseteq\mathbb{X}^{2}_{t}$ for any $t$. In filtration on $\mathbb{X}^{1}$, the $1$-cycle $\tau$ is terminated by $\sigma_{j}$ at a death time $f_{1}(\sigma_{j})$. That is, $\tau$ is zero homologous in $\mathbb{X}^{1}_{f_{1}(\sigma_{j})}$. However, since $\mathbb{X}^{1}_{f_{1}(\sigma_{j})}\subseteq\mathbb{X}^{2}_{f_{1}(\sigma_{j})}$, $\tau$ is also terminated at $f_{1}(\sigma_{j})$ in filtration $\mathbb{X}^{2}$. Therefore, we obtain $f_{2}(\sigma_{k})\leq f_{1}(\sigma_{j})$ because the value $f_{2}(\sigma_{k})$ is a death time of $\tau$, which means the first time to die in filtration $\mathbb{X}^{2}$.

### v2

@@ 3.2 Toy example of path-representable distance
Consider a connected weighted graph $G=(V,E,W_{E})$ with $V=\{v_{1},v_{2},v_{3},v_{4}\}$ and $E=$ $\{(v_{1},v_{2}),$ $(v_{2},v_{3})$, $(v_{3},v_{4})$$(v_{1},v_{4})\}$ as shown in the left figure of Fig. 3. How many path-representable distances can we make for any path choice function $g$ on $G$? A maximal path can represent any path choice function $g$ on $G$, where the maximal path refers to a path that is maximal with respect to the inclusion. The right figure of Fig. 3 represents the maximal paths indicated by the given path choice function on graph $G$. The maximal paths are well-defined due to the consistency property of the path choice function. In this case, a total of eight possible path choice functions can be defined on $G$, as shown in Fig. 3.
@@
For the connected weighted graph $G=(V,E,W_{E})$ shown in the left figure of Fig. 4, all path-representable distances derived from all the path choice functions are illustrated in the right figure of Fig. 4, with a total of eight possibilities. In this case, all the distances are weight-dominated, while only the distances $d_{5},d_{6},d_{7},d_{8}$ are cost-dominated.
@@ 3.3 Main theorem
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider two cost-dominated path-representable distances $d_{i}$ and $d_{j}$ by path choice functions $g_{i}$ and $g_{j}$ on $G$, respectively. If $d_{i}(v,w)\leq d_{j}(v,w)$ for any vertices $v$ and $w$, then there exists an injective function $\varphi_{i,j}:\textbf{bcd}_{1}(d_{i})\rightarrow\textbf{bcd}_{1}(d_{j})$ defined as $\varphi_{i,j}([\beta,\delta])=[\beta,\delta^{\prime}]$ such that $\delta\leq\delta^{\prime}$ for any $\beta,\delta]\in\textbf{bcd}_{1}(d_{i})$.
Figure 5: The injective functions $\varphi_{i,j}:\textbf{bcd}_{1}(d_{i})\rightarrow\textbf{bcd}_{1}(d_{j})$, as described in Theorem 3.4, for the four cost-dominated path-representable distances $d_{5},d_{6},d_{7},d_{8}$ in Fig. 4.
Fig. 5 illustrates the application of Theorem 3.4 to the cost-dominated path-representable distances $d_{5},d_{6},d_{7},d_{8}$ described in Fig. 4. The left figure of Fig. 5 shows the partial order relationships among distances $d_{5},d_{6},d_{7},d_{8}$, while the right figure displays the injective functions $\varphi_{5,6},\varphi_{6,7},\varphi_{5,8},\varphi_{8,7}$ between their 1-dimensional persistence barcodes.
@@ 3.4 Remarks
The main theorem can also be formulated using the concept of a poset. Let $G$ be a connected weighted graph. To analyze the graph by defining the distance between the nodes, it is natural to define the set whose elements are the distances defined on $G$. Especially in our theory, we focus on the poset $(\mathcal{L},\leq)$ whose elements are the cost-dominated path-representable distances defined on $G$, where for two elements $d_{1},d_{2}\in\mathcal{L}$, $d_{1}\leq d_{2}$ if $d_{1}(v,w)\leq d_{2}(v,w)$ for any two points $v,w\in V$. Furthermore, we calculate persistent homology using the standard algorithm [17], which calculates a persistence barcode from the given distance matrix. Note that $(G,d)$ gives a unique distance matrix. We define such a function as $\textbf{bcd}_{k}$, that is, $\textbf{bcd}_{k}(d)$ is a $k$-dimensional persistence barcode obtained by $(G,d)$. Then the codomain of the function will be the poset of persistence barcodes on $G$, as a subset of the poset of interval sets paired with $k$-simplices. Here, the partial order holds if we can find the proper inclusion between the elements with the same associated simplex in the sense of the main theorem. To be precise, for two $k$-dimensional persistence barcodes $S_{1}$ and $S_{2}$, we define $S_{1}\leq S_{2}$ if there is an injection $\varphi:S_{1}\hookrightarrow S_{2}$ such that $\varphi([\beta,\delta]_{\sigma})=[\beta,\delta^{\prime}]_{\sigma}\text{ for any }[\beta,\delta]_{\sigma}\in S_{1}\text{ with }\delta\leq\delta^{\prime}.$ The main theorem claims that $\textbf{bcd}_{1}$ becomes the order-preserving map.
@@
First, if we choose $d_{\text{weight}}$ as an element, then it becomes the least element according to Definition 2.1. Note that $d_{\text{weight}}$ is indeed the element in the domain by Proposition 3.3.
Now we take a counterexample for the greatest element. Let $G$ be a cycle graph with at least 5 vertices and suppose that $d_{\max}$ is the greatest element in $P$. As illustrated in Fig. 6, fix a control point $v_{0}$. Since $d_{\max}$ is path-representable, for every other vertex $v$ of $G$, we must choose between two paths from $v_{0}$ to $v$ which $d_{\max}$ calculates. Let $v_{i}$ and $v_{i+1}$ be two consecutive nodes. Let $d_{1}$ be the distance that chooses the path from $v_{0}$ to $v_{i}$ that does not contain $v_{i+1}$ and the path from $v_{0}$ to $v_{i+1}$ that contains $v_{i}$. Let $d_{2}$ be the distance that chooses the path from $v_{0}$ to $v_{i+1}$ that does not contain $v_{i}$ and the path from $v_{0}$ to $v_{i}$ that contains $v_{i+1}$. Note that we can make these two distances cost-dominated since $v_{i}$ and $v_{i+1}$ are not adjacent to $v_{0}$, and $G$ is a cycle graph. Then $d_{1}(v_{0},v_{i})>d_{2}(v_{0},v_{i})$ and $d_{1}(v_{0},v_{i+1})<d_{2}(v_{0},v_{i+1})$. Thus, due to the assumption for $d_{\max}$, we have $d_{\max}(v_{0},v_{i})=d_{1}(v_{0},v_{i})$ and $d_{\max}(v_{0},v_{i+1})=d_{2}(v_{0},v_{i+1})$ since there are only two candidates for each selection of path. This leads to a contradiction to the consistency condition of path-representable distance. ∎
Figure 7: $\textbf{bcd}_{1}$ being an order preserving map between the poset of cost-dominated path-representable distance on $G$ (Left) and the poset of 11-dimensional persistence barcode on $G$ from Fig. 4 (Right). Here the arrow denotes the partial order for each poset.
@@ 4 Proofs of main theory
For a weighted graph $G=(V,E,W_{E})$, let $d:V\times V\rightarrow\mathbb{R}$ be any distance defined on $G$. If $\sigma=(v,w)$ is a birth simplex of dimension $1$ with the distance $d(v,w)$ of a $1$-dimensional cycle, there is no vertex $v_{0}\in V(G)$ such that $d(v,v_{0})<d(v,w)$ and $d(w,v_{0})<d(v,w)$.
@@
Let $e=(v,w)$ be a birth simplex of dimension $1$ with a birth value $d(v,w)$. Suppose $e\notin E$. There is a path $g(v,w)$ in $G$ such that $d(v,w)=\sum_{e\in g(v,w)}W_{E}(e)$. Denote the path $g(v,w)$ as the sequence of vertices $v=v_{0},v_{1},v_{2},\ldots,v_{m-1},v_{m}=w$ and $a_{k}=W_{E}((v_{k-1},v_{k}))$. Then, by the consistency property of the path choice function $g$, the path $g(v,w_{m-1})$ is a sequence of $v_{0}$, $v_{1}$, $v_{2}$, $\ldots$, $v_{m-1}$. Therefore, we know that $d(v,v_{m-1})=a_{1}+a_{2}+\cdots+a_{m-1}<d(v,w)$. Since $d(v_{m-1},w)<d(v,w)$ is obvious, this contradicts Lemma 4.1. ∎
Let $\mathcal{B}$ be the set of all birth edges in $\textbf{bcd}_{1}(d)$. By Proposition 4.2, we know that $\mathcal{B}\subseteq E$. Generally, for any edge $e=(v,w)\in E$, it is not necessary that $W_{E}(e)=d(v,w)$. For example, cases where $W_{E}(e)$ $d_{\text{weight}}(v,w)$ are common. However, for the birth edges, the following interesting proposition holds.
@@
Let $G=(V,E,W_{E})$ be any connected weighted graph and $\mathcal{T}\in\text{MST}(G)$. Suppose that we have a cycle $\tau$ in $G$ and that there is an edge $e\in\tau$ such that $W_{E}(e)>W_{E}(e^{\prime})$ for any other edge $e^{\prime}\in\tau$. Then $e\notin\mathcal{T}$.
@@
Consider a connected weighted graph $G=(V,E,W_{E})$. Define $\bar{E}=\{(v,w)\mid v,w\in V\}$. For any path choice function $g$, we have its path-representable distance $d_{g}$ and a weighted complete graph $K^{g}=(V,\bar{E},W_{\bar{E}})$, where $W_{\bar{E}}(e)=d_{g}(v,w)$ for all $e=(v,w)\in\bar{E}$. We name this graph $K^{g}$ the graph completion of $G$ by $g$.
@@
(iii) Consider a minimum spanning tree $\mathcal{T}^{\prime}\in\text{MST}(G)$. If $\mathcal{T}=\mathcal{T}^{\prime}$, then the proposition holds. Suppose that $\mathcal{T}\neq\mathcal{T}^{\prime}$, which means that there is an edge $e\in\mathcal{T}\setminus\mathcal{T}^{\prime}$ and a unique path $p^{\prime}$ in $\mathcal{T}^{\prime}$ connecting the endpoints of $e$. In the path $p^{\prime}$, there is an edge $e^{\prime}\in\mathcal{T}^{\prime}\setminus\mathcal{T}$ because $e+p^{\prime}$ is a cycle in $G$ and $\mathcal{T}$ is a tree. By $(ii)$, we have $W_{E}(e)=W_{\bar{E}}(e)$. Suppose that $W_{E}(e^{\prime})\neq W_{\bar{E}}(e^{\prime})$ and denote $e^{\prime}=(v,w)$. Then there is a path $g(v,w)$ such that $W_{E}(e^{\prime})>W_{E}(e_{i})$ for any $e_{i}\in g(v,w)$. This contradicts Lemma 4.4 because $e^{\prime}\in\mathcal{T}^{\prime}\in\text{MST}(G)$ so that $W_{E}(e^{\prime})=W_{\bar{E}}(e^{\prime})$. Since $\mathcal{T}^{\prime}\in\text{MST}(G)$, we have $W_{E}(e^{\prime})\leq W_{E}(e)$. Otherwise, by replacing $e^{\prime}$ with $e$ in the edge swap process, we have a contradiction as $\mathcal{T}^{\prime}\in\text{MST}(G)$. Similarly, since $\mathcal{T}\in\text{MST}(K^{g})$, we have $W_{\bar{E}}(e^{\prime})\geq W_{\bar{E}}(e)$. To summarize, the following hold:
@@
Proposition 4.6 implies that any weight-dominated path choice function $g$ always keeps $\text{MST}(G)$ invariant, for which the weight-dominated condition is essential. Otherwise, neither $\text{MST}(G)$ nor $\text{MST}(K^{g})$ can be guaranteed to be larger. A counterexample is shown in Fig. 8 below.
@@
For the fixed order $\pi$ on the edge set $\bar{E}$ in $K^{g}$, consider the set of all $1$-simplices $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$ for some 0-simplex $\sigma_{i}$, denoted by $\mathcal{T}^{\pi}$. Then
[…]
@@
$(\supseteq)$ The process of sequentially constructing a minimum spanning tree matches precisely the process of incrementally calculating $\text{low}_{\mathbf{R}}$. To explain in more detail, for every vertex in $K^{g}$, each 0-simplex (vertex) is connected to a $1$-simplex (edge) causing its death. According to the definition of $\text{low}_{\mathbf{R}}$, no cycles can form hence it is a spanning tree. Since $\text{low}_{\mathbf{R}}$ selects edges in order from the lowest value, it becomes the minimum spanning tree. Therefore, for any order $\pi$ on the edge set $\bar{E}$ in $K^{g}$, $\mathcal{T}^{\pi}$ is an element in $\text{MST}(K^{g})$.
$(\subseteq)$ Consider an arbitrary element $\mathcal{T}\in\text{MST}(K^{g})$. Let $\pi$ be an order on $\bar{E}$ such that, for any given weights on the edges, if the weights are the same, the edges in $\mathcal{T}$ are preferred over those in $K^{g}\setminus\mathcal{T}$. When computing the persistent homology of $K^{g}$, even if the order of the $1$-simplices (edges) corresponding to the indices of the matrix in the $\mathbf{R}=\mathbf{D}\mathbf{V}$ algorithm is rearranged according to $\pi$, the result of the persistent homology computation remains unchanged (permutation invariant). In this case, $\mathcal{T}=\mathcal{T}^{\pi}$. ∎
@@
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider any weight-dominated choice functions $g_{1}$ and $g_{2}$ and their completions $K^{g_{1}}$ and $K^{g_{2}}$. Then there exists an order $\pi$ in $\bar{E}$ such that their minimum spanning trees on $K^{g_{1}}$ and $K^{g_{2}}$ induced by the standard algorithm are exactly the same.
@@
(i) First, we show that the $e$-column of the reduced $R_{1}$ is zero. From $e\in\mathcal{B}_{2}$, we know that the $e$-column of $\mathbf{R}_{2}$ is zero. The fact that the $e$-column of $\mathbf{R}_{2}$ is a zero column means that there are column indices $j_{1}$, $\ldots$, $j_{m}$ corresponding to edges $e_{j_{1}},\ldots,e_{j_{m}}$ such that $e+\sum_{l=1}^{m}e_{j_{l}}$ is a cycle and there are 0-simplices $\sigma_{i_{j_{k}}}$ such that $i_{j_{k}}=\text{low}_{\mathbf{R}_{2}}(j_{k})$. By Proposition 4.8, we can assume that the minimum spanning trees $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ for $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$ in the standard algorithm are exactly the same for some fixed order $\pi$ on edge set $\bar{E}$. Therefore, we know that $i_{j_{k}}=\text{low}_{\mathbf{R}_{1}}(j_{k})$ for all $k$. Thus the $e$-column of $\mathbf{R}_{1}$ is zero due to the column indices $j_{1},\ldots,j_{m}$ during the column operations in the standard algorithm.
(ii) Next, we show that if the $1$-cycle $\tau$ corresponding to the birth edge $e=\sigma_{i}$ is terminated by the $2$-simplex $\sigma_{j}$ in $\mathbf{R}_{1}$, then $f_{1}(\sigma_{i})<f_{1}(\sigma_{j})$. Consider a pair of simplices $(\sigma_{i},\sigma_{j})$ which is $\text{low}_{\mathbf{R}_{1}}(j)=i$. For any $t\in\mathbb{R}$, define $\mathbb{X}^{2}_{t}={f_{2}}^{-1}((-\infty,t])$ and $\mathbb{X}^{1}_{t}={f_{1}}^{-1}((-\infty,t])$. As $f_{2}\leq f_{1}$ on $\mathbb{X}$, we know that $\mathbb{X}^{1}_{t}\subseteq\mathbb{X}^{2}_{t}$ for any $t$. In filtration on $\mathbb{X}^{1}$, the $1$-cycle $\tau$ is terminated by $\sigma_{j}$ at a death value $f_{1}(\sigma_{j})$. That is, $\tau$ is zero homologous in $\mathbb{X}^{1}_{f_{1}(\sigma_{j})}$. However, since $\mathbb{X}^{1}_{f_{1}(\sigma_{j})}\subseteq\mathbb{X}^{2}_{f_{1}(\sigma_{j})}$, $\tau$ is also terminated at $f_{1}(\sigma_{j})$ in filtration $\mathbb{X}^{2}$. Therefore, we obtain $f_{2}(\sigma_{k})\leq f_{1}(\sigma_{j})$ because the value $f_{2}(\sigma_{k})$ is a death value of $\tau$, which means the first value to die in filtration $\mathbb{X}^{2}$.